OpenStudy (hxkage):

If PQR is an equilateral triangle and PX is perpendicular to QR, then PR^2 is equal to: a) 3/2 RX^2 b) 2 RX^2 c) 3 RX^2 d) 4 RX^2

OpenStudy (hxkage):



OpenStudy (visarika23):

PQR is an equilateral triangle therefore perpendicular PX will bisect the side QR into two equal halves. Let PQ=QR=PR=a Then RX=a/2 PR^2=a^2 RX^2=a^2/4 So PR^2=4RX^2

OpenStudy (spider_mastermind_4):

If the question is answered please close the question @hxkage

OpenStudy (hxkage):

@Visarika23 Thank you! ;3

OpenStudy (hxkage):

@Spider_Mastermind_4 Sorry, didn't have connection for the last two days. ;-;

OpenStudy (visarika23):

Wlcm☺